OpenStudy (anonymous):

fully factor? (2x+1)^2 - (3x-2)^2

OpenStudy (jdoe0001):

\(\bf (2x+1)^2 - (3x-2)^2\implies (2x+1)(2x+1) - (3x-2)(3x-2)\)

OpenStudy (anonymous):

i got there, and then i got this: 2x^2+4x+1-9x^2-12x+4

OpenStudy (anonymous):

which simplifies to: -7x^2-8x+5, but it doesn't factor nicely from there...?

OpenStudy (jdoe0001):

ohhh shoot.. you said factor... I see

OpenStudy (jdoe0001):

my bad... one sec... recall that \(\bf \large a^2-b^2 = (a-b)(a+b)\)

OpenStudy (jdoe0001):

what you were doing was simplifying, but thtey just want it factored

OpenStudy (anonymous):

yeah i put it into standard form so i could factor it but it doesn't factor nicely?

OpenStudy (jdoe0001):

firstly \(\bf (2x+1)^2 - (3x-2)^2\implies [4x^2+4x+1]-[9x^2-12x+4]\\ \quad \\ 4x^2+4x+1-9x^2+12x-4\implies -5x^2+16x-3\)

OpenStudy (jdoe0001):

secondly, I gather that once you get the "standard form", then you'd factor it, you'd end up with a form like \(\bf a^2-b^2 = (a-b)(a+b)\) anyway

OpenStudy (anonymous):

oops i guess i should know how to add if i'm doing math... i got it now thanks!